find the remaining parts of the quadrantal spherical triangles (c=90°) given the following, giveb parts b= 45°20 a=76°40

1
Expert's answer
2022-02-22T08:29:06-0500

Cos"90°=\frac{Cos\>C-Cos45°20'Cos76°40'}{Sin45°20'\>Sin76°40'}"


"C=80\u00b040'"


"\frac{Sin\>A°}{Sin76°40'}=\frac{Sin90°}{Sin80°40'}"


"A=80\u00b026'"


"\frac{SinB°}{Sin\>45°20'}=\frac{Sin\>90°}{Sin\>80°40'}"


"B=46\u00b07'"